OpenStudy (anonymous):

selby walked for a while at 4mph and jogged the rest of the way at 8 mph. if she covered 56 miles in 10 hours how far did she walk and how far did she jog?

OpenStudy (zarkon):

solve \[\frac{x}{4}+\frac{y}{8}=10\] \[x+y=56\] x=walking miles and y=jogging miles

OpenStudy (anonymous):

let the distance walked at rate of 4 mph = x miles => time taken to cover x miles = t1 = x/4 h also distance jogged at 8 mph = (56 - x) miles => time taken to cover (56 - x) miles = t2 = (56 - x)/8 h total time = t1 + t2 so x 56 - x 10 = ------ + ------- 4 8 solve this for x and you will get yr answer....

OpenStudy (anonymous):

x 56 - x 10 = ------ + ------- 4 8 80 = 2x + 56 - x 80 - 56 = x 24 = x

OpenStudy (anonymous):

willyg - see the question about buses